The bottom of Jim's rectangular bait box is 3 inches longer than it is wide. The diagonal is 15 inches. What is the area of the bottom of the box???
Answer by htmentor(1343)

The bottom of Jim's rectangular bait box is 3 inches longer than it is wide. The diagonal is 15 inches. What is the area of the bottom of the box???
======================
Let w = the width of the box
Then the length = w + 3
Since the diagonal is 15, we have a right triangle with sides w, w+3 and hypotenuse 15
Solve using the Pythagorean theorem:
w^2 + (w+3)^2 = 15^2
This simplifies to
w^2 + 3w - 108 = 0
which can be factored as:
(w+12)(w-9) = 0
Take the positive solution: w = 9
So the area = 9*12 = 108 sq. in.